# Score Distillation Sampling with Learned Manifold Corrective

### v1

@@ 4 Method
Our proposal is to compare $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z},t)$ and $\boldsymbol{\epsilon}$ in $\mathcal{L}_{\text{proj}}$, the predicted denoised image $\hat{\mathbf{z}}_{t}=\mathbf{z}_{\boldsymbol{\phi}}(\mathbf{z},t)$, with the rendered image $\mathbf{z}$. To this end, we rewrite $\mathcal{L}_{\text{proj}}$ as
$\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}=\left(\frac{\sqrt{\alpha_{t}}}{\sqrt{1-\alpha_{t}}}\left(\mathbf{z}-\hat{\mathbf{z}}_{t}\right)\right)\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}},$ (7)
which is equivalent to comparing the predicted and true noise. By looking at Eq. 7, we can now see a possible explanation why $\mathcal{L}_{\text{proj}}$ might not be optimal: $\hat{\mathbf{z}}_{t}$ is an approximation of $\mathbf{z}$ and most probably not a perfect replica of $\mathbf{z}$, even when $\mathbf{z}$ already lies on the natural image manifold, see Fig. 1. This is especially the case for large values of $t$, where the denoising model has to reconstruct the true image from almost pure noise. In an attempt to better understand the behavior of the denoising model given $\mathbf{z}$ and $t$, we model it as a two step process, as follows
$\mathbf{z}_{\boldsymbol{\phi}}(\mathbf{z},t)=b\circ p(\mathbf{z},t)=b(\tilde{\mathbf{z}},t),$ (8)
@@
where $p$ projects $\mathbf{z}$ onto the learned natural image manifold, $\tilde{\mathbf{z}}$ is the projected image, and $\tilde{\mathbf{z}}=\mathbf{z}$ for images on the manifold. $b$ is the error or bias of the denoising model introduced by the current time step $t$. In this two-step model we are only interested in $p$ and would like to neglect any gradients coming from $b$. As it is unfeasible to correct for $b$, we propose to approximate $b$ with $\hat{b}$ and apply it on $\mathbf{z}$ before comparing it with $\hat{\mathbf{z}}_{t}$
@@
$\small\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{lmc}}=\left(\hat{b}(\mathbf{z},t)-\hat{\mathbf{z}}_{t}\right)\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}}.$ (9)
@@
In this formulation we omit both the Jacobian term from the denoising model (like in the original formulation, cf. Eq. 3) as well as the Jacobian from $\hat{b}$. Also, we empirically found that by dropping the weighting term $\frac{\sqrt{\alpha_{t}}}{\sqrt{1-\alpha_{t}}}$ we can further improve visual fidelity – which is similar to the common practice of predicting the original natural image instead of the noise when learning the denoising model. As evident in Fig. 1, given an adequate model $\hat{b}$, $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{lmc}}$ focuses on global changes, whereas $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}$ is dominated by high-frequent detail. We now explain how to obtain $\hat{b}$.
@@
We propose to model $\hat{b}$ as the learnable function $\hat{b}_{\boldsymbol{\psi}}(\mathbf{z},t)$ parameterized by $\boldsymbol{\psi}$. We can observe $b$ as a function of $t$ by sampling triplets $(\mathbf{z},\hat{\mathbf{z}}_{t},t)$, where $\mathbf{z}$ are random natural images assumed to be approximately on manifold images, i.e. $\mathbf{z}\approx\tilde{\mathbf{z}}$. Using these triplets we can learn $\hat{b}_{\boldsymbol{\psi}}$ by minimizing $\|\hat{\mathbf{z}}_{t}-\hat{b}_{\mathbf{z},t}\|^{2}_{2}$, with $\hat{b}_{\mathbf{z},t}=\hat{b}_{\boldsymbol{\psi}}(\mathbf{z},t)$. Using a non-deterministic model, $\hat{b}_{\boldsymbol{\psi}}$ is expected to under-fit the highly complex and probabilistic nature of $b$. In fact, for a given pair $(\mathbf{z},t)$ at best it can learn the mean over all possible $\hat{\mathbf{z}}_{t}$. This is desired, because even after approximating $b(\mathbf{z},t)$ through $\hat{b}_{\boldsymbol{\psi}}(\mathbf{z},t)$, we will be able to observe a gradient in $\mathcal{L}_{\text{lmc}}$: a gradient towards a specific instance of $\hat{\mathbf{z}}_{t}$. Further, $\hat{b}_{\boldsymbol{\psi}}$ learns to “blur” images and is not informed by the image manifold associated to the diffusion model. It will perform similar operations for on-manifold and off-manifold images and thus $\mathcal{L}_{\text{lmc}}$ may correct non-manifold images. However, naively applying a $\hat{b}_{\boldsymbol{\psi}}$ in Eq. 9, learned that way, is causing an undesired effect, namely a drift in image statistics. Concretely we observed that images produced by $\hat{b}_{\boldsymbol{\psi}}$ tend to underestimate the true image dynamics, especially for large $t$ where $\hat{\mathbf{z}}_{t}$ may be vastly different to $\mathbf{z}$. By underestimating image dynamics, however, $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{lmc}}$ contains a gradient to “correct” for that and we have found images obtained via optimization using this gradient to be overly saturated and contrasting. To this end, we make $\hat{b}_{\boldsymbol{\psi}}$ agnostic to the global image statistics by computing the training loss in a normalized space
@@
The original SDS formulation has been reported to have mode-seeking behaviour. Consequently, optimization results obtained via $\mathcal{L}_{\text{SDS}}$ have the tendency to look very similar, even when based on different random seeds. We hypothesize this is partly due to optimizing over many steps and hence averaging the effects of varying noise $\boldsymbol{\epsilon}$. To prevent such averaging behaviour, we propose to fix $\boldsymbol{\epsilon}$ over the course of optimization – which is in spirit similar to DDIM sampling [35]. Empirically, we found that optimizing with fixed $\boldsymbol{\epsilon}$ is often unstable in the original SDS formulation. We hypothesize this is rooted in $\mathcal{L}_{\text{proj}}$ not providing proper gradients towards the image manifold and at the same time $\mathcal{L}_{\text{cond}}$ providing too strong gradients towards prompt agreement. The optimization state may leave the manifold of natural images and drift towards a state associated more to an “adversarial example” for the given prompt. While our loss formulation allows the use of a fixed $\boldsymbol{\epsilon}$, probably due to both the usage of smaller $\omega$ and better manifold correctives, we observe that we can further robustify the formulation by only fixing $\boldsymbol{\epsilon}$ in $\mathcal{L}_{\text{cond}}$, while continue to sample $\boldsymbol{\epsilon}$ in $\mathcal{L}_{\text{lmc}}$. This results in high quality, yet diverse, results, cf. Sec. 5.
@@
We use an Imagen version [33] trained on internal data sources as the pre-trained text-to-image diffusion model $\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}$ or $\mathbf{z}^{\omega}_{\boldsymbol{\phi}}$, respecively. We use only the base model and not the super-resolution cascade, producing images of $128\times 128$ pixel resolution. We model $\hat{b}_{\boldsymbol{\psi}}$ as a standard U-Net with four encoder layers with two Conv/ReLU/MaxPool blocks per layer, and skip connections to the decoder. The first layer contains 32 filters and the number of filters is doubled in each layer in the encoder and halved in the decoder. $\hat{b}_{\boldsymbol{\psi}}$ has been trained until convergence, using $\mathcal{L}_{\text{k}}$ as the only loss. For creating training triplets $(\mathbf{z},\hat{\mathbf{z}}_{t},t)$, we sampled random images $\mathbf{z}$ from OpenImages [17], random time-steps $t\sim\mathcal{U}(0.02,0.98)$, and passed both into $\mathbf{z}_{\boldsymbol{\phi}}$ obtaining $\hat{\mathbf{z}}_{t}$.
@@ 5 Experiments
Figure 2: Optimization-based image synthesis results. We optimize an empty image to match the given prompt using our LMC-SDS, the original SDS formulation, and $\mathcal{L}_{\text{cond}}$. SDS struggles to create detailed content when using low guidance $\omega$. High $\omega$ produces detailed results but colors may be oversaturated (chimpanzee face), fake detail may appear (2nd mouse tail), or artifacts emerge. $\mathcal{L}_{\text{cond}}$ is unstable to optimize and produces unrealistic colors. In contrast, our method produces detailed results with balanced colors.
@@
Set-up.
If not noted otherwise, we scale the total loss by $\frac{1}{\omega}$ to be able to ablate $\omega$ without changing the gradient magnitude. The pre-trained diffusion model we use in our experiments produces images of $128\times 128$px resolution. We enable arbitrary higher resolutions in our image experiments through the following strategy: We downsample the optimization state to $128\times 128$px for the first iterations. This ensures that the whole image is affected by the loss, but the signal may be blurry. Accounting for this, we sample random patches of arbitrary resolutions, from the image, for the remaining optimization steps. To prevent the loss from producing fake detail in sampled image patches, we lower the learning rate during this phase by using Cosine decay.
Baselines.
This paper presents a generic novel loss function which uses an image diffusion model. To this end, we compare with the original SDS formulation and with related work focused on improving SDS. Specifically, we compare against DDS [8] and multi-step SDS (MS-SDS) [39]. DDS and MS-SDS originally were proposed using a latent diffusion model [30] and operate in latent space. For fair comparisons, we adapted both methods to operate in pixel space and to use the same image diffusion model. Note that optimizing in the (ambient) pixel space is generally considered harder due higher-dimensionality. Finally, we evaluate the effectiveness of our LMC-SDS loss for 3D asset generation using DreamFusion. Hereby we focus on the improvements of our loss and deliberately exclude recent improvements over DreamFusion rooted in other factors, e.g. novel-view synthesis networks or additional perceptual losses. While being important findings, those factors are orthogonal and complementary to our approach.
Metrics.
We use the following metrics for evaluating our methods and baselines: We report Learned Perceptual Image Patch Similarity (LPIPS $\downarrow$) [38] w.r.t. the source image, thus capturing how much an image was altered. Note that high LPIPS can be a result of either successful editing or image corruption. To disambiguate these, we therefore additionally report the CLIP score ($\uparrow$) [9] w.r.t. the target prompt. As the CLIP score can be insensitive, we additionally report CLIP-R-Precision ($\uparrow$) [27], the retrieval accuracy using the CLIP model, where feasible.
@@ 5.1 Image Synthesis & Editing
Following the patching strategy described earlier, we can create images of arbitrary resolutions. To this end, we initialize an image with 50% grey and optimize its pixel values to match a given prompt. We show results of our loss and SDS in Fig. 2. We additionally include results obtained by only using $\mathcal{L}_{\text{cond}}$, ablating the influence of our proposed $\mathcal{L}_{\text{lmc}}$. Optimizing solely with $\mathcal{L}_{\text{cond}}$ is unstable and the final images feature clearly visible defects, such as oversaturated colors or noisy patterns. The prompt is also not always well respected as e.g. in the case of the gibbon. The original SDS formulation produces blurry and desaturated images when using low guidance $\omega$. High $\omega$ leads to better performance but artifacts are present, e.g. fake detail like the second mouse tail, oversaturated colors like in the chimpanzee’s face, or blocky artifacts. In contrast, our method produces images of higher visual fidelity which respect the prompt well.
@@
Our loss can also be used for optimization-based image editing. We load existing natural images and optimize their pixel values directly. Optionally, we may constrain the optimization with an L2 loss w.r.t. the original pixel values. In Fig. 3 we show qualitative results of so edited images along with the corresponding target prompt. All results faithfully respect the target prompt while keeping the image structure intact. We can also vary the result by fixing $\boldsymbol{\epsilon}$ over the course of the optimization, cf. Sec. 4. In Fig. 4, we show an input image along with four result variants. All results respect the target prompt well, while at the same time being significantly different from each other.
@@
Figure 5: Quantitative results for optimization-based editing under varying $\omega$. DDS stays close to the original image (low LPIPS) by performing only small edits (low CLIP). SDS and MS-SDS respect the prompt better (higher CLIP), but corrupt the image (high LPIPS). Our method shows the strongest editing capabilities (highest CLIP), while staying close to the original structure.
Figure 5: Quantitative results for optimization-based editing under varying $\omega$. DDS stays close to the original image (low LPIPS) by performing only small edits (low CLIP). SDS and MS-SDS respect the prompt better (higher CLIP), but corrupt the image (high LPIPS). Our method shows the strongest editing capabilities (highest CLIP), while staying close to the original structure.
@@
Figure 6: Comparison of image statistics before (Original) and after editing with $\omega=15$. We plot the average power spectra and vertical derivative histograms. Ours and DDS preserve the image statistics well, while SDS-based methods introduce significant blur.
@@
Figure 7: Results obtained with cats-to-others image translation networks trained with different losses. The network trained with the SDS loss fails to create detailed results. The DDS network produces rather uniform results, whereas the network trained with our LMC-SDS creates realistic and detailed images.
For a numerical evaluation, we use an internal dataset of 1,000 images with text editing instructions in form of source and target prompts, similar to the dataset used in InstructPix2Pix [2]. We optimize images to match the target prompt solely using our LMC-SDS and baselines DDS, SDS, and MS-SDS. For MS-SDS we choose to use five step denoising. We also experimented with more steps, but didn’t observe significantly different performance. We report numerical results with varying $\omega$ (without loss rescaling) in Fig. 6. We found DDS to be least effective and producing results closest to the original image. We hypothesize that the negative gradient w.r.t. to the original image may not always be helpful or even turn out to be “confusing”. This could also explain a behavior reported by original authors, where editing results on real images are “inferior” over those on synthetic images. A possible reason being, that since for synthetic images the original prompts are describing the image well, the negative gradients are less “confusing”. SDS can effectively edit the image, but also corrupts the image. This is also evident from the altered image statistics, see Fig. 6. In our experiments, MS-SDS performs very similar to SDS, corrupting the image in similar manner, but also being slightly less effective for editing. MS-SDS produces clear and detailed denoised images as part of the loss computation. We hypothesize that since denoised images and the current optimization state may differ drastically, and as denoised images vary considerably at each step, the computed gradients are not always very meaningful. The original paper proposes MS-SDS to be used with a view-conditioned diffusion model, where the denoised images probably vary much less and computed gradients may be more expedient. Our loss respects the prompt best (highest CLIP), and finds a good balance between preserving the original structure and performing significant edits.

### v2

@@ 4 Method
We now present our solution to eliminating the frequency bias in $\mathcal{L}_{\text{proj}}$ that we identified earlier. In an attempt to better understand the behavior of the denoising model given $\mathbf{z}$ and $t$, we model it as a two step process, as follows
$\mathbf{z}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t)=b\circ p(\mathbf{z}_{t},t)=b(\tilde{\mathbf{z}},t),$ (8)
@@
where $p$ projects $\mathbf{z}$ onto the learned natural image manifold, $\tilde{\mathbf{z}}$ is the projected image, and $\tilde{\mathbf{z}}=\mathbf{z}$ for images on the manifold. $b$ is the error or frequency bias of the denoising model introduced by the current time step $t$. In this two-step model we are only interested in $p$ and would like to neglect any gradients coming from $b$. As it is unfeasible to correct for $b$, we propose to cancel out effects of $b$ by approximating $b$ with $\hat{b}$ and applying $\hat{b}$ on $\mathbf{z}$ before comparing $\mathbf{z}$ with $\hat{\mathbf{z}}_{t}$:
@@
$\small\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{lmc}}=\left(\hat{b}(\mathbf{z},t)-\hat{\mathbf{z}}_{t}\right)\frac{\partial\mathbf{z}_{t}}{\partial\boldsymbol{\theta}}$ (9)
@@
In this formulation we omit both the Jacobian term from the denoising model (like in the original formulation, cf. Eq. 3) as well as the Jacobian from $\hat{b}$. Also, we empirically found that by dropping the weighting term $\frac{\sqrt{\alpha_{t}}}{\sqrt{1-\alpha_{t}}}$ we can further improve visual fidelity – which is similar to the common practice of predicting the original image instead of the noise when learning the denoising model. As demonstrated in Fig. 1 (left), given an adequate model $\hat{b}$, $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{lmc}}$ focuses on global changes, whereas $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{proj}}$ is dominated by high-frequent detail. We now explain how to obtain $\hat{b}$.
@@
We propose to model $\hat{b}$ as a learnable function $\hat{b}_{\boldsymbol{\psi}}(\mathbf{z},t)$ parameterized by $\boldsymbol{\psi}$. We can observe $b$ as a function of $t$ by sampling triplets $(\mathbf{z},\hat{\mathbf{z}}_{t},t)$, where $\mathbf{z}$ are random natural images assumed to be approximately on manifold images, i.e. $\mathbf{z}\approx\tilde{\mathbf{z}}$. Using these triplets we can learn $\hat{b}_{\boldsymbol{\psi}}$ by minimizing $\|\hat{\mathbf{z}}_{t}-\hat{b}_{\mathbf{z},t}\|^{2}_{2}$, with $\hat{b}_{\mathbf{z},t}=\hat{b}_{\boldsymbol{\psi}}(\mathbf{z},t)$. Modeled with a deterministic model, $\hat{b}_{\boldsymbol{\psi}}$ is expected to under-fit the highly complex and probabilistic nature of $b$. In fact, for a given pair $(\mathbf{z},t)$ at best it can learn the mean over all possible $\hat{\mathbf{z}}_{t}$. This is desired, because even after approximating $b(\mathbf{z},t)$ through $\hat{b}_{\boldsymbol{\psi}}(\mathbf{z},t)$, we will be able to observe a gradient in $\mathcal{L}_{\text{lmc}}$: a gradient towards a specific instance of $\hat{\mathbf{z}}_{t}$. Further, $\hat{b}_{\boldsymbol{\psi}}$ only learns to “blur” images and neither generates new content nor is informed by the diffusion models’ image manifold. Thus, it will perform similar operations for on-manifold and off-manifold images and $\mathcal{L}_{\text{lmc}}$ may correct non-manifold images. However, naively applying a $\hat{b}_{\boldsymbol{\psi}}$ in Eq. 9, learned that way, is causing an undesired effect, namely a drift in image statistics. Concretely we observed that images produced by $\hat{b}_{\boldsymbol{\psi}}$ tend to underestimate the true image dynamics, especially for large $t$ where $\hat{\mathbf{z}}_{t}$ may be vastly different to $\mathbf{z}$. By underestimating image dynamics, however, $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{lmc}}$ contains a gradient to “correct” for that and we have found images obtained via optimization using this gradient to be overly saturated and contrasting. To this end, we make $\hat{b}_{\boldsymbol{\psi}}$ agnostic to the global image statistics by computing the training loss in a normalized space
@@
The original SDS formulation has been reported to have mode-seeking behaviour. Consequently, optimization results obtained via $\mathcal{L}_{\text{SDS}}$ have the tendency to look very similar, even when based on different random seeds. We hypothesize this is due to optimizing over many steps and hence averaging the effects of varying noise $\boldsymbol{\epsilon}$. To prevent such averaging behaviour, we propose to fix $\boldsymbol{\epsilon}$ over the course of optimization – which is in spirit similar to DDIM sampling [36]. Empirically, we found that optimizing with fixed $\boldsymbol{\epsilon}$ is often unstable in the original SDS formulation. We hypothesize this is rooted in $\mathcal{L}_{\text{proj}}$ not providing proper gradients towards the image manifold and at the same time $\mathcal{L}_{\text{cond}}$ providing too strong gradients towards prompt agreement. The optimization state may leave the manifold of natural images and drift towards a state associated more to an “adversarial example” for the given prompt. While our loss formulation allows the use of a fixed $\boldsymbol{\epsilon}$, probably due to both the usage of smaller $\omega$ and better manifold correctives, we observe that we can further robustify the formulation by only fixing $\boldsymbol{\epsilon}$ in $\mathcal{L}_{\text{cond}}$, while continue to sample $\boldsymbol{\epsilon}$ in $\mathcal{L}_{\text{lmc}}$. This results in high quality, yet diverse, results, cf. Sec. 5.
@@
We use an image diffusion model producing images of $128\times 128$ pixel resolution and trained on internal data sources [34] as the pre-trained text-to-image diffusion model $\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\phi}}$ or $\mathbf{z}^{\omega}_{\boldsymbol{\phi}}$, respecively. We model $\hat{b}_{\boldsymbol{\psi}}$ as a standard U-Net with four encoder layers with two Conv/ReLU/MaxPool blocks per layer, and skip connections to the decoder. The first layer contains 32 filters and the number of filters is doubled in each layer in the encoder and halved in the decoder. $\hat{b}_{\boldsymbol{\psi}}$ has been trained until convergence, using $\mathcal{L}_{\text{k}}$ as the only loss. For creating training triplets $(\mathbf{z},\hat{\mathbf{z}}_{t},t)$, we sampled random images $\mathbf{z}$ from OpenImages [18], random time-steps $t\sim\mathcal{U}(0.02,0.98)$, and passed both into $\mathbf{z}_{\boldsymbol{\phi}}$ obtaining $\hat{\mathbf{z}}_{t}$. We train $\hat{b}_{\boldsymbol{\psi}}$ once and keep it fixed for all experiments.
@@ 5 Experiments
Figure 2: Optimization-based image synthesis results. We optimize an empty image to match a given prompt using our LMC-SDS, the original SDS loss, and $\mathcal{L}_{\text{cond}}$. SDS struggles to create detailed content when using low guidance $\omega$. High $\omega$ produces detailed results but colors may be oversaturated (chimpanzee face), fake detail may appear (2nd mouse tail), or artifacts emerge. $\mathcal{L}_{\text{cond}}$ is unstable to optimize and produces unrealistic colors. In contrast, our method produces detailed results with balanced colors.
Set-up.
If not noted otherwise, we scale the total loss by $\frac{1}{\omega}$ to be able to ablate $\omega$ without changing the gradient magnitude. Fixed $\boldsymbol{\epsilon}$ sampling is only used when explicitly stated. The pre-trained diffusion model we use in our experiments produces images of $128\times 128$px resolution. We enable arbitrary higher resolutions in our image experiments through the following strategy: We downsample the optimization state to $128\times 128$px for the first iterations. This ensures that the whole image is affected by the loss, but the signal may be blurry. Accounting for this, we sample random patches of arbitrary resolutions, from the image, for the remaining optimization steps. To prevent the loss from producing fake detail in sampled image patches, we lower the learning rate during this phase through Cosine decay.
Baselines.
This paper presents a generic novel loss function which uses an image diffusion model. To this end, we compare with the original SDS formulation and with related work focused on improving SDS. Specifically, we compare against DDS [8], NFSD [13], and multi-step SDS (MS-SDS) [40]. DDS, NFSD, and MS-SDS originally were proposed using a latent diffusion model [31] and operate in latent space. For fair comparisons, we adapted all methods to operate in pixel space and to use the same image diffusion model. Note that optimizing in the (ambient) pixel space is generally considered harder due to higher-dimensionality. Finally, we evaluate the effectiveness of our LMC-SDS loss for 3D asset generation using DreamFusion. Hereby we focus on the improvements of our loss and deliberately exclude recent improvements over DreamFusion rooted in other factors, e.g. novel-view synthesis networks or additional perceptual losses. While being important findings, those factors are orthogonal and complementary to our approach.
Metrics.
We use the following metrics for evaluating our methods and baselines: We report Learned Perceptual Image Patch Similarity (LPIPS $\downarrow$) [39] w.r.t. the source image, thus capturing how much an image was altered. Note that high LPIPS can be a result of either successful editing or image corruption. To disambiguate these, we additionally report the CLIP score ($\uparrow$) [9] w.r.t. the target prompt. As CLIP can be insensitive, we additionally report CLIP-R-Precision ($\uparrow$) [28], the retrieval accuracy using the CLIP model, where feasible.
@@ 5.1 Image Synthesis & Editing
To be able to evaluate and ablate the original SDS loss and our LMC-SDS in isolation (without additional interfering losses), we propose the task of optimization-based image synthesis. Instead of using the denoising schedule of a diffusion model, we initialize an image with 50% grey and optimize its pixel values to match a given prompt, following the patching strategy described earlier. While it is not expected that images obtained using this strategy are on par with those obtained via the generation process, the results allow to draw meaningful conclusions about the properties of the loss functions. We show results of our loss and SDS in Fig. 2. We additionally include results obtained by only using $\mathcal{L}_{\text{cond}}$, ablating the influence of our proposed $\mathcal{L}_{\text{lmc}}$. Optimizing solely with $\mathcal{L}_{\text{cond}}$ is unstable and the final images feature clearly visible defects, such as oversaturated colors or noisy patterns. The prompt is also not always well respected as e.g. in the case of the gibbon. The original SDS formulation produces blurry and desaturated images when using low guidance $\omega$. High $\omega$ leads to better performance but artifacts are present, e.g. fake detail like the second mouse tail, oversaturated colors like in the chimpanzee’s face, or blocky artifacts. In contrast, our method produces images of higher visual fidelity which respect the prompt well. Additionally, we empirically found that problems of the individual losses become more pronounced when optimizing longer (here we only optimize for 500 steps, cf. Sec. 5.2 for effects after 50K steps).
@@
Our loss can also be used for optimization-based image editing. We load existing natural images and optimize their pixel values directly. Optionally, we may constrain the optimization with an L2 loss w.r.t. the original pixel values. In Fig. 3 (more in Supp. Mat.) we show qualitative results of so edited images along with the corresponding target prompt. All results faithfully respect the target prompt while keeping the image structure intact. We can also vary the result by fixing $\boldsymbol{\epsilon}$ over the course of the optimization, cf. Sec. 4. In Fig. 4, we show an input image along with four result variants. All results respect the target prompt well, while at the same time being significantly different from each other.
Refer to captionRefer to caption
@@
Figure 5: Quantitative results for optimization-based editing under varying $\omega$. Left: Our method results in the highest CLIP score over all baselines for all $\omega$. Right: We plot LPIPS over CLIP for further performance insights: DDS stays close to the original image (lowest LPIPS) by performing only small edits (low CLIP). SDS and MS-SDS respect the prompt better (higher CLIP), but corrupt the image (high LPIPS). NFSD corrupts the image less (lower LPIPS), but exhibits weak editing capabilities (low CLIP). Our method shows the strongest editing capabilities (highest CLIP), while staying close to the original structure.
@@
Figure 6: Comparison of image statistics before (Original) and after editing with $\omega=5$. We plot the average power spectra (left) and vertical derivative histograms (right). Ours, DDS, and NFSD preserve the image statistics well, while SDS-based methods introduce significant blur.
@@
Figure 7: Results obtained with cats-to-others image translation networks trained with different losses. The network trained with SDS and NFSD fail to create detailed results. The DDS network produces rather uniform results, whereas the network trained with our LMC-SDS creates realistic and detailed images.
@@
For a numerical evaluation, we use an internal dataset of 1,000 images with text editing instructions in form of source and target prompts, similar to the dataset used in InstructPix2Pix [2]. We optimize images to match the target prompt solely using our LMC-SDS and baselines DDS, NFSD, SDS, and MS-SDS. For MS-SDS we chose to use five step denoising. We also experimented with more steps, but didn’t observe significantly different performance. We report numerical results with varying $\omega$ (without loss rescaling) in Fig. 6. We found DDS to be least effective and producing results closest to the original image. We hypothesize that the negative gradient w.r.t. to the original image may not always be helpful. This could also explain a behavior reported by original authors, where editing results on real images are “inferior” over those on synthetic images. A possible reason being, that since for synthetic images the original prompts are describing the image well, the negative gradients are more helpful in that case. SDS can effectively edit the image, but also corrupts the image. This is also evident from the altered image statistics, see Fig. 6. NFSD corrupts the image less, but exhibits weak editing capabilities. In our experiments, MS-SDS performs very similar to SDS, corrupting the image in similar manner, but also being slightly less effective for editing. MS-SDS produces clear and detailed denoised images as part of the loss computation. We hypothesize that since denoised images and the current optimization state may differ drastically, and as denoised images vary considerably at each step, the computed gradients are not always very meaningful. The original paper proposes MS-SDS to be used with a view-conditioned diffusion model, where the denoised images probably vary much less and computed gradients may be more directed. Our loss respects the prompt best (highest CLIP), and finds a good balance between preserving the original structure and performing significant edits.
To quantify the how much the result diversity can be increased through fixing $\boldsymbol{\epsilon}$, we re-ran the experiment using LMC-SDS ($\omega=30$) ten times with different seeds, both using the standard strategy and with fixing $\boldsymbol{\epsilon}$ in $\mathcal{L}_{\text{cond}}$. We then computed the LPIPS (lower is more similar) of nine runs w.r.t. to the 10th (reference) run. The results using the standard strategy are all very similar with a per-run average of $0.041\pm 0.007$ w.r.t. the reference run. In contrast, our new strategy produces much more diverse results: $0.158\pm 0.018$.